# Central limit theorems for interacting innovation processes, related statistical tools and general results

### v3

@@ 3. Statistical inference
We here present some statistical tools based on the results of Section 2. More precisely, we provide a statistical test for each of the two interaction matrices, $\Gamma$ and $W$, based on the observable processes ${\mathbf{D}}^{*}_{t}$ and ${\mathbf{K}}_{t}(c)$. Furthermore, we use the latter process to provide also a confidence interval for the limit probability $\widetilde{P}_{\infty}(c)$. The general procedure to be followed in order to obtain these statistical instruments is described in Sections B C. In particular, in our framework, we have to apply it with ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$ (or, equivalently ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}$) and with ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$. To this end, we observe that Assumption B.1 is true in both cases with $\Sigma_{det}=\Sigma_{det,\Gamma}=diag({\mathbf{u}})$ in the first one and $\Sigma_{det}=\Sigma_{det,W}=I$ in the second one.
We firstly focus on the case $N=2$ and, then, we consider a particular type of interaction (mean-field interaction) in the general case $N\geq 2$. It is worthwhile to underline that the statistical tools described in Section B are very general and so it is also possible to explicit the formulas for any family of interaction matrices.
@@ 3.1.2. Hypothesis test on $W$
and so we can make inference on the two parameters $w_{2,1}$ and $w_{1,2}$. Using the parametration $\iota=w_{1,2}+w_{2,1}$ for the intensity of the interaction due to the matrix $W$, and $\eta=\tfrac{w_{1,2}}{w_{1,2}+w_{2,1}}$ (resp. $(1-\eta)$) for the percentage of this intensity due to the influence of process $1$ on process $2$ (resp., of process $2$ on process $1$), we have
@@ 3.1.3. Confidence interval for $\widetilde{P}_{\infty}(c)$
As an example, we have simulated $S=200$ independent pair of innovation processes with $\Gamma$ and $W$ defined with $r=0.75$, $\gamma^{*}=0.75$, $\eta=1/2$, $\iota_{\Gamma}=1$ and $\iota_{W}=1.25$ and then, for any simulation, we have constructed the confidence intervals (3.9) of level $(1-\alpha)=95\%$ computed at time-step $t=10^{3}$ associated to the item $c$ with highest $(K_{t,1}(c)+K_{t,2}(c))$ (recall that the interval has the same expression for any given item $c$). In Figure 1 we can see these $S=200$ confidence intervals (sorted according to the value of $K_{t,1}(c)+K_{t,2}(c)\,$) and the estimates of the corresponding $\widetilde{P}_{\infty}(c)$ computed as $\tfrac{K_{t_{\infty},1}(c)+K_{t_{\infty},2}(c)}{2t_{\infty}}$ at time-step $t_{\infty}=10^{5}$. The empirical coverage is represented by the proportion of times the estimates of $\widetilde{P}_{\infty}(c)$ belong to the corresponing confidence interval, and in this simulation this is equal to $0.955$, which is very close to the nominal value of $0.95$.
Figure 1. Confidence interval (see (3.9)) for the random variable $\widetilde{P}_{\infty}(c)$ at time-step $t=10^{3}$ and related to item $c$ with highest ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$ in the case $N=2$, $\Gamma$ defined as in (3.1) with $r=0.75$, $\gamma^{*}=0.75$, $\eta=1/2$ and $\iota_{\Gamma}=1$, and $W$ defined as in (3.5) with $\eta=1/2$ and $\iota_{W}=1.25$. Number of simulations $S=200$ (sorted according to the value of ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$). The crosses represent the value of $\widetilde{P}_{\infty}(c)$ estimated by ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t_{\infty}}(c)/(Nt_{\infty})$ at time-step $t_{\infty}=10^{5}$.
@@ 3.2.3. Confidence interval for $\widetilde{P}_{\infty}(c)$
Figure 2. Confidence interval (see (3.14)) for the random variable $\widetilde{P}_{\infty}(c)$ at time-step $t=10^{3}$ and related to item $c$ with highest ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$ in the case $N=3$, $\Gamma$ defined as in (3.10) with $\phi=0.75$ and $\iota_{\Gamma}=0.8$, and $W$ defined as in (3.12) with $\iota_{W}=0.8$. Number of simulations $S=200$ (sorted according to the value of ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$). The crosses represent the value of $\widetilde{P}_{\infty}(c)$ estimated by ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t_{\infty}}(c)/(Nt_{\infty})$ at time-step $t_{\infty}=10^{5}$.
@@ 4. Applications
In this section we apply the proposed inferential methodology on the two real data sets, already considered in Aletti et al. (2023), to test the minimum level of interaction presented by the matrices $\Gamma$ and $W$ in these specific contexts. In particular, the first one is taken from the social content aggregation website Reddit, collected, elaborated and made freely available on the web by the authors of Monti et al. (2023) (see https://github.com/corradomonti/demographic-homophily), while the second one is downloaded from the on-line library Project Gutenberg, which is a collection of public domain books (see https://www.gutenberg.org).
@@ 4.1. Reddit data set
This data set consists of a collection of news, and comments associated to each news, for the period 201620202016-2020, downloaded from the r/news community on the website Reddit Monti et al. (2023). Each news is associated with the author who posted it. Moreover, the data set contains the specific topic the news belongs to and, to each comment is also assigned a measurement of the sentiment, expressed as a real value in (1,1)(-1,1). We consider only the comments to news belonging to the topic ”Politics“. Moreover, we categorize the sentiment variable: precisely, we define it as ”positive“ if the provided sentiment value was larger than +0.35+0.35 and ”negative“ if the provided sentiment value was lower than 0.35-0.35. (We remove any comment with an original sentiment value that lies within (0.35,+0.35)(-0.35,+0.35).) We finally take a data set of 2 602 1732\,602\,173 comments for each of the two categories. Hence, we consider two innovation processes: one related to the authors who receive at least one comment with negative sentiment for the news they post and the similar one referred to the comments with positive sentiment. A first analysis of this data set can be found in Aletti et al. (2023), where we illustrate that these two sequences exhibit empirical behaviours that are in accordance with those predicted by Theorems 1.1 and 1.2, and we provide the estimates $\widehat{\gamma}^{*}=0.781$ and $\widehat{r}=10^{-0.727}=0.187$ (these estimates have been obtained by means of a suitable regression on $({\mathbf{D}}^{*}_{t})_{t}$, which simulations have shown to be a highly accurate estimation method). We here continue the study performing the two hypothesis tests on $\Gamma$ (taking the above estimates as the known values of $\gamma^{*}$ and $r$) and on $W$ described in Subsection 3.1. More precisely, we consider the two one-sided tests:
@@
$\iota_{\Gamma,0}$ p-value $\iota_{W,0}$ p-value: $q_{50\%}$,  $q_{75\%}$,  $q_{95\%}$
@@ 4.2. Gutenberg data set
This data set comes from the on-line library Project Gutenberg and it is related to seven books written in English and belonging to two particular literary genres: “Western” and “’History”. Precisely, it consists of 476 948476\,948 words for both genres. The two sequences of words are the innovation processes object of the study. Therefore, each word is new or old for a specific literary genre if, respectively, it has or has not already been used within that genre. Analogously, each word is new or old for the entire system if, respectively, it has or has not already been used within any considered genre. A first analysis of also this data set can be found in Aletti et al. (2023), where we show that the behaviors of these two sequences predicted by Theorems 1.1 and 1.2 match with the observed ones, and we provide the estimates $\widehat{\gamma}^{*}=0.466$ and $\widehat{r}=10^{-0.238}=0.578$. As in the previous application, we here continue the study performing the two hypothesis tests on $\Gamma$ and $W$ described in Subsection 3.1. More precisely, we consider the two one-sided tests:
@@
$\iota_{\Gamma,0}$ p-value $\iota_{W,0}$ p-value: $q_{50\%}$,  $q_{75\%}$,  $q_{95\%}$
@@
In all the appendix, the notation $R_{t}=O(s_{t})$ means that $|R_{t}|\leq Cs_{t}$ for a suitable constant $C$ and for $t$ large enough. Similarly, the vectorial notation ${\mathbf{R}}_{t}=O(s_{t}){\boldsymbol{1}}$ means that each component of ${\mathbf{R}}_{t}$, say $R_{t,j}$, is such that $|R_{t,j}|\leq Cs_{t}$ for a suitable constant $C$ and for $t$ large enough.
@@ Appendix A General results
    $\Phi^{\top}$ is a non-negative irreducible matrix with leading eigenvalue $0<\phi^{*}\leq 1$;
@@
    $\Delta{\mathbf{M}}_{t+1}=O(t^{1-\eta}){\boldsymbol{1}}$ with $\eta>\phi^{*}/2$ and $t^{-(1-\phi^{*})}E[\Delta{{\mathbf{M}}}_{t+1}\Delta{{\mathbf{M}}}_{t+1}^{\top}|\mathcal{F}_{t}]\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}\Sigma_{\infty}$ with $\Sigma_{\infty}$ a positive semidefinite symmetric random matrix.
It is worthwhile to note that dynamics similar to (A.1) have already been considered in the Stochastic Approximation literature (e.g. Laruelle and Pagès (2013); Mokkadem and Pelletier (2006)). However, this kind of results do not apply to our situation. Indeed, the covariance matrices in those results are deterministic, while the asymptotic covariance matrices in our CLTs are random. This is also why we do not use the simple convergence in distribution, but we employ the notion of stable convergence, which is, among other things, essential for obtaining the statistical tools.
@@
Theorem A.1.
(CLT for $(\widetilde{A}_{t})$ when $N\geq 1$)
In addition to assumptions $(i)$, $(ii)$ and $(iii)$, also assume:
@@
    $\widetilde{A}_{t}\stackrel{{\scriptstyle a.s./in\,mean}}{{\longrightarrow}}\widetilde{A}_{\infty}$, where $\widetilde{A}_{\infty}$ is an integrable random variable.
Then, we have
$t^{\phi^{*}/2}(\widetilde{A}_{t}-\widetilde{A}_{\infty})\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}(0,\phi^{*}{\mathbf{v}}^{\top}\Sigma_{\infty}{\mathbf{v}})\,.$
Moreover, the above convergence is in the sense of the almost sure conditional convergence with respect to the filtration $(\mathcal{F}_{t})$.
@@
Since $\widetilde{A}_{t}\stackrel{{\scriptstyle a.s./in\,mean}}{{\to}}\widetilde{A}_{\infty}$ and $\sum_{n\geq 0}\widetilde{R}_{A,n+1}=\sum_{n\geq 0}O(1/n^{1+\beta})<+\infty$  then $\breve{M}_{t}\stackrel{{\scriptstyle a.s./in\,mean}}{{\to}}\breve{M}_{\infty}$ for a certain integrable random variable $\breve{M}_{\infty}$. Moreover, we have
@@
where $\beta-\phi^{*}/2>0$. Summing up it is enough to prove the convergence of $t^{\phi^{*}/2}(\breve{M}_{t}-\breve{M}_{\infty})$. For this purpose, we want to apply Theorem D.1. Let us consider, for each $t$ the filtration $({\mathcal{F}}_{t,k})_{k}$ and the process $(L_{t,k})_{k}$ defined by
@@
where the almost sure convergence follows by applying (Crimaldi et al., 2016, Lemma 4.1) and by the fact that
@@
Theorem A.2.
(Convergence result for $(\widehat{A}_{t},\widehat{B}_{t})$ when $N=1$)
@@
Theorem A.3.
(CLT for $(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})$ when $N\geq 2$)
In addition to assumptions $(i)$, $(ii)$ and $(iii)$, let $\phi_{2}^{*}$ be an eigenvalue of $\Phi$ different from $\phi^{*}$ with highest real part, i.e. $\phi_{2}^{*}\in Sp(\Phi)\setminus\{\phi^{*}\}$ such that $\mathcal{R}e(\phi_{2}^{*})=\max\{\mathcal{R}e(\phi):\,\phi\in Sp(\Phi)\setminus\{\phi^{*}\}\}$ and assume
@@
where, denoting by $\phi_{j}$ the eigenvalues of $\Phi$ different from $\phi^{*}$, we have for $j=2,\dots,N$
@@
Firstly, we recall that the matrix $\Phi^{\top}$ can be decomposed as
$\Phi^{\top}=\phi^{*}{\mathbf{u}}{\mathbf{v}}^{\top}+UDV^{\top}\,$,
where $D$ is the diagonal matrix whose elements are the eigenvalues of $\Phi$ (i.e. of $\Phi^{\top}$) different from $\phi^{*}$ and $U$ and $V$ denote the matrices whose columns are the left (right) and the right (left) eigenvectors of $\Phi$ (of $\Phi^{\top}$, respectively) associated to these eigenvalues, so that we have
Then, we have $\widehat{{\mathbf{A}}}_{t}=UV^{\top}{\mathbf{A}}_{t}$ and the dynamics of $\widehat{{\mathbf{A}}}_{t}$ can be obtained by multypling by $UV^{\top}$ the dynamics of ${\mathbf{A}}_{t}$ as follows:
@@
Denoting the eigenvalues of $\Phi^{\top}$ different from $\phi^{*}$ by $\phi_{j}$ for $j=2,\dots,N$, let us set $\alpha_{j}=\phi^{*}-\phi_{j}$. Recall that $\mathcal{R}e(\alpha_{j})>0$ since $\phi^{*}$ is the leading eigenvalue and so $\mathcal{R}e(\phi_{j})<\phi^{*}$. Then, setting an initial time $m_{0}$ such that ${\mathcal{R}e}(\alpha_{j})/(m_{0}+1)<1$ for each $j$, we can write
@@
(A.5) $\begin{split}C_{k+1,t}&=U_{\boldsymbol{\eta}}A_{k+1,t}V_{\boldsymbol{\eta}}^{\top},\qquad\mbox{and}\\ A_{k+1,t}&=\begin{cases}\prod_{m=k+1}^{t}\left(I-\frac{1}{m+1}S_{Q}\right)=\begin{pmatrix}A^{11}_{k+1,t}&{\boldsymbol{0}}&0\\ 0&a^{22}_{k+1,t}&0\\ A^{31}_{k+1,t}&{\boldsymbol{0}}&A^{33}_{k+1,t}\\ \end{pmatrix}\quad&k=m_{0}-1,\dots,t-1,\\ I\quad&k=t\,.\end{cases}\end{split}$
Notice that the blocks $A^{11}_{k+1,t}$, $A^{31}_{k+1,t}$ and $A^{33}_{k+1,t}$ are all diagonal $(N-1)\times(N-1)$ matrices. Moreover, setting for any $x\in\mathbb{C}$ with ${\mathcal{R}e}(x)/(m_{0}+1)<1$, $p_{m_{0}-1}(x)=1$ and $p_{k}(x)=\prod_{m=m_{0}}^{k}(1-\tfrac{x}{m+1})$ for $k\geq m_{0}$ and $F_{k+1,t}(x)=\frac{p_{t}(x)}{p_{k}(x)}$ for $m_{0}-1\leq k\leq t-1$, we get (see the supplementary material of Aletti et al. (2019) for similar technical calculations)
@@
To this purpose, setting $B_{{\boldsymbol{\eta}},k+1}=V_{\boldsymbol{\eta}}^{\top}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}^{\top}V_{\boldsymbol{\eta}}$, $B_{k+1}=V^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}V$, ${\mathbf{b}}_{k+1}=V^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}{\mathbf{v}}$ and $b_{k}={\mathbf{v}}^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}{\mathbf{v}}$, we observe that
@@
Morever, with some computations (see the supplementary material of Aletti et al. (2020) for similar technical calculations) for complex numbers $x$, $y$ such that ${\mathcal{R}e}(x+y)-\phi^{*}>0$ we have

### v4

@@ 4. Statistical inference
We here present some statistical tools based on the results of Section 3. More precisely, we provide a statistical test for each of the two interaction matrices, $\Gamma$ and $W$, based on the observable processes ${\mathbf{D}}^{*}_{t}$ and ${\mathbf{K}}_{t}(c)$. Furthermore, we use the latter process to provide also a confidence interval for the limit probability $\widetilde{P}_{\infty}(c)$. The general procedure to be followed in order to obtain these statistical instruments is described in Section B and Section C. In particular, in our framework, we have to apply it with ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$ (or, equivalently ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}$) and with ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$. To this end, we observe that Assumption B.1 is true in both cases with $\Sigma_{det}=\Sigma_{det,\Gamma}=diag({\mathbf{u}})$ in the first one and $\Sigma_{det}=\Sigma_{det,W}=I$ in the second one.
We firstly focus on the case $N=2$ and, then, we consider a particular type of interaction (mean-field interaction) in the general case $N\geq 2$. It is worthwhile to underline that the statistical tools described in Section B are very general and so it is also possible to explicit the formulas for any family of interaction matrices satisfying the required assumptions.
@@ 4.1.2. Hypothesis test on $W$
and so we can make inference on the two parameters $w_{2,1}$ and $w_{1,2}$. Using the parametrization $\iota=w_{1,2}+w_{2,1}$ for the intensity of the interaction due to the matrix $W$, and $\eta=\tfrac{w_{1,2}}{w_{1,2}+w_{2,1}}$ (resp. $(1-\eta)$) for the percentage of this intensity due to the influence of process $1$ on process $2$ (resp., of process $2$ on process $1$), we have
@@ 4.1.3. Confidence interval for $\widetilde{P}_{\infty}(c)$
As an example, we have simulated $S=200$ independent pair of innovation processes with $\Gamma$ and $W$ defined with $r=0.75$, $\gamma^{*}=0.75$, $\eta=1/2$, $\iota_{\Gamma}=1$ and $\iota_{W}=1.25$ and then, for any simulation, we have constructed the confidence intervals (4.9) of level $(1-\alpha)=95\%$ computed at time-step $t=10^{3}$ associated to the item $c$ with highest $(K_{t,1}(c)+K_{t,2}(c))$ (recall that the interval has the same expression for any given item $c$). In Figure 1 we can see these $S=200$ confidence intervals (sorted according to the value of $K_{t,1}(c)+K_{t,2}(c)\,$) and the estimates of the corresponding $\widetilde{P}_{\infty}(c)$ computed as $\tfrac{K_{t_{\infty},1}(c)+K_{t_{\infty},2}(c)}{2t_{\infty}}$ at time-step $t_{\infty}=10^{5}$. The empirical coverage is represented by the proportion of times the estimates of $\widetilde{P}_{\infty}(c)$ belong to the corresponding confidence interval, and in this simulation this is equal to $0.955$, which is very close to the nominal value of $0.95$.
Figure 1. Confidence interval (see (4.9)) for the random variable $\widetilde{P}_{\infty}(c)$ at time-step $t=10^{3}$ and related to item $c$ with highest ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$ in the case $N=2$, $\Gamma$ defined as in (4.1) with $r=0.75$, $\gamma^{*}=0.75$, $\eta=1/2$ and $\iota_{\Gamma}=1$, and $W$ defined as in (4.5) with $\eta=1/2$ and $\iota_{W}=1.25$. Number of simulations $S=200$ (sorted on the $x$-axis according to the value of ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$). The crosses represent the value (reported on the $y$-axis) of $\widetilde{P}_{\infty}(c)$ estimated by ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t_{\infty}}(c)/(Nt_{\infty})$ at time-step $t_{\infty}=10^{5}$.
@@ 4.2.3. Confidence interval for $\widetilde{P}_{\infty}(c)$
Figure 2. Confidence interval (see (4.14)) for the random variable $\widetilde{P}_{\infty}(c)$ at time-step $t=10^{3}$ and related to item $c$ with highest ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$ in the case $N=3$, $\Gamma$ defined as in (4.10) with $\phi=0.75$ and $\iota_{\Gamma}=0.8$, and $W$ defined as in (4.12) with $\iota_{W}=0.8$. Number of simulations $S=200$ (sorted on the $x$-axis according to the value of ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t}(c)$). The crosses represent the value (reported on the $y$-axis) of $\widetilde{P}_{\infty}(c)$ estimated by ${\boldsymbol{1}}^{\top}{\mathbf{K}}_{t_{\infty}}(c)/(Nt_{\infty})$ at time-step $t_{\infty}=10^{5}$.
@@ 5. Applications
In this section we apply the proposed inferential methodology on the two real data sets, already considered in Aletti et al., (2023), to test the minimum level of interaction presented by the matrices $\Gamma$ and $W$ in these specific contexts. In particular, the first one is taken from the social content aggregation website Reddit, collected, elaborated and made freely available on the web by the authors of Monti et al., (2023) (see https://github.com/corradomonti/demographic-homophily), while the second one is downloaded from the on-line library Project Gutenberg, which is a collection of public domain books (see https://www.gutenberg.org).
@@ 5.1. Reddit data set
The data set includes a collection of news posts and their corresponding comments from the Reddit community r/news over the period 2016201620202020, downloaded from the website Reddit Monti et al., (2023). Each news item is linked to the user who posted it. In addition, the data provide the topic associated with each news post and assign to every comment a sentiment score, expressed as a real number in the interval (1,1)(-1,1). We restrict our analysis to comments on news classified under the topic “Politics”. The sentiment variable is then discretized as follows: a comment is labeled as “positive” if its sentiment score exceeds +0.35+0.35, and as “negative” if it is below 0.35-0.35. Comments with original sentiment values in (0.35,+0.35)(-0.35,+0.35) are excluded. The resulting data set contains 2 602 1732\,602\,173 comments in each of the two categories. Consequently, we define two innovation processes: one associated with authors receiving at least one negative comment on their posts, and another associated with authors receiving at least one positive comment. A first analysis of this data set can be found in Aletti et al., (2023), where we illustrate that these two sequences exhibit empirical behaviours that are in accordance with those predicted by Theorems 2.1 and 2.2, and we provide the estimates $\widehat{\gamma}^{*}=0.781$ and $\widehat{r}=10^{-0.727}=0.187$ (these estimates have been obtained by means of a suitable regression on $({\mathbf{D}}^{*}_{t})_{t}$, which simulations have shown to be a highly accurate estimation method). We here continue the study performing the two hypothesis tests on $\Gamma$ (taking the above estimates as the known values of $\gamma^{*}$ and $r$) and on $W$ described in Subsection 4.1. More precisely, we consider the two one-sided tests:
@@
Interaction Intensity $\iota_{\Gamma,0}$ p-value Interaction Intensity $\iota_{W,0}$ p-value distribution $q_{50\%}$, $q_{75\%}$, $q_{95\%}$
@@ 5.2. Gutenberg data set
This data set is drawn from the on-line library Project Gutenberg and concerns seven books written in English that belong to two specific literary genres: “Western” and “History”. More precisely, the corpus contains 476 948476\,948 words for each genre. The two resulting word sequences constitute the innovation processes under investigation. A word is classified as new (respectively, old) for the overall system if it has not (respectively, has already) appeared in any of the considered genres. A first analysis of also this data set can be found in Aletti et al., (2023), where we show that the behaviors of these two sequences predicted by Theorems 2.1 and 2.2 match with the observed ones, and we provide the estimates $\widehat{\gamma}^{*}=0.466$ and $\widehat{r}=10^{-0.238}=0.578$. As in the previous application, we here continue the study performing the two hypothesis tests on $\Gamma$ and $W$ described in Subsection 4.1. More precisely, we consider the two one-sided tests:
@@
Interaction Intensity $\iota_{\Gamma,0}$ p-value Interaction Intensity $\iota_{W,0}$ p-value distribution $q_{50\%}$, $q_{75\%}$, $q_{95\%}$
@@
In all the appendix, the notation $R_{t}=O(s_{t})$ means that $|R_{t}|\leq Cs_{t}$ for a suitable deterministic constant $C$ and for $t$ large enough. Similarly, the vectorial notation ${\mathbf{R}}_{t}=O(s_{t}){\boldsymbol{1}}$ means that each component of ${\mathbf{R}}_{t}$, say $R_{t,j}$, is such that $|R_{t,j}|\leq Cs_{t}$ for an appropriate deterministic constant $C$ and for $t$ large enough. (Analogous results can be obtained for a suitable integrable random variable $C$.)
The results in appendix are given in terms of stable convergence or its variants. For more details on these concepts, we refer the reader to Aldous and Eagleson, (1978); Crimaldi, (2016); Crimaldi et al., (2007); Hall and Heyde, (1980); Rényi, (1963).
@@ Appendix A General results
    $\Phi^{\top}$ is a non-negative irreducible diagonalizable matrix with leading eigenvalue $0<\phi^{*}\leq 1$;
@@
    $\Delta{\mathbf{M}}_{t+1}=O(t^{1-\eta}){\boldsymbol{1}}$ is a martingale difference with $\eta>\phi^{*}/2$ and $t^{-(1-\phi^{*})}E[\Delta{{\mathbf{M}}}_{t+1}\Delta{{\mathbf{M}}}_{t+1}^{\top}|\mathcal{F}_{t}]\stackrel{{\scriptstyle a.s.}}{{\rightarrow}}\Sigma_{\infty}$ with $\Sigma_{\infty}$ a positive semidefinite symmetric random matrix.
It is worthwhile to note that dynamics similar to (A.1) have already been considered in the Stochastic Approximation literature. However, the most part of those results do not apply to our situation because the covariance matrices in those results are deterministic (e.g. Laruelle and Pagès, (2013); Mokkadem and Pelletier, (2006)), while the asymptotic covariance matrices in our CLTs are random. This is also why we do not use the simple convergence in distribution, but we employ the notion of stable convergence. On the other hand, the paper Zhang, (2016) also takes into account random asymptotic covariance matrices, but it is able to catch our CLT only in the case $\phi^{*}=1$ (compare our assumption $(iii)$ by assumption (2.4) in Zhang, (2016)).
@@
An intuitive reasoning of this decomposition relates to the eigen-structure of the matrix $(\phi^{*}I-\Phi^{\top})$ in (A.1). Indeed, $\widetilde{A}_{t}$ is the projection along its kernel, that corresponds to the Perron Frobenius eigen-space of $\Phi$, and $\widehat{{\mathbf{A}}}_{t}$ completes with the projections on the remaining directions. The process $({\mathbf{B}}_{t})_{t}$ in (A.1) will follow the dynamics of the attracting process $({\mathbf{A}}_{t})_{t}$, and $\widehat{{\mathbf{B}}}_{t}$ plays the role of $\widehat{{\mathbf{A}}}_{t}$.
Theorem A.1 (CLT for $(\widetilde{A}_{t})$ when $N\geq 1$).
In addition to assumptions $(i)$, $(ii)$ and $(iii)$, also assume
@@
    $(\widetilde{A}_{t})_{t}$ is uniformly integrable.
Then $(\widetilde{A}_{t})_{t}$ converges almost surely and in mean to an integrable random variable $\widetilde{A}_{\infty}$ and we have
$t^{\phi^{*}/2}(\widetilde{A}_{t}-\widetilde{A}_{\infty})\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}(0,\phi^{*}{\mathbf{v}}^{\top}\Sigma_{\infty}{\mathbf{v}})\,$,
Moreover, the above convergence is in the sense of the almost sure conditional convergence with respect to the filtration $(\mathcal{F}_{t})_{t}$.
@@
Since assumption $(iv)$ and $\sum_{n\geq 0}\widetilde{R}_{A,n+1}=\sum_{n\geq 0}O(1/n^{1+\beta})<+\infty$, we obtain that $\breve{M}_{t}\stackrel{{\scriptstyle a.s./in\,mean}}{{\to}}\breve{M}_{\infty}$ for a certain integrable random variable $\breve{M}_{\infty}$ and so $\widetilde{A}_{t}\stackrel{{\scriptstyle a.s./in\,mean}}{{\to}}\widetilde{A}_{\infty}$. Moreover, we have
@@
where $\beta-\phi^{*}/2>0$. Therefore, it is enough to prove the convergence of $t^{\phi^{*}/2}(\breve{M}_{t}-\breve{M}_{\infty})$. For this purpose, we want to apply Theorem D.1. Let us consider, for each $t$ the filtration $({\mathcal{F}}_{t,k})_{k}$ and the process $(L_{t,k})_{k}$ defined by
@@
where the almost sure convergence follows by applying (Crimaldi et al.,, 2016, Lemma 4.1) and by the fact that
@@
Theorem A.2 (Convergence result for $(\widehat{A}_{t},\widehat{B}_{t})$ when $N=1$).
@@
We recall that the diagonalizable and irreducible matrix $\Phi^{\top}$ can be decomposed as
(A.3) $\Phi^{\top}=\phi^{*}{\mathbf{u}}{\mathbf{v}}^{\top}+UDV^{\top}\,$,
where $D$ is the diagonal matrix whose elements are the eigenvalues $\phi_{j}$, $j\geq 2$ of $\Phi$ (i.e. of $\Phi^{\top}$) different from $\phi^{*}$ and $U$ and $V$ denote the matrices whose columns are the left (right) and the right (left) eigenvectors of $\Phi$ (of $\Phi^{\top}$, respectively) associated to these eigenvalues, such that we have
Let $\phi_{2}^{*}$ be an eigenvalue of $\Phi$ different from $\phi^{*}$ with highest real part, i.e. $\phi_{2}^{*}\in Sp(\Phi)\setminus\{\phi^{*}\}$ such that $\mathcal{R}e(\phi_{2}^{*})=\max\{\mathcal{R}e(\phi):\,\phi\in Sp(\Phi)\setminus\{\phi^{*}\}\}$.
Theorem A.3 (CLT for $(\widehat{{\mathbf{A}}}_{t},\widehat{{\mathbf{B}}}_{t})$ when $N\geq 2$).
In addition to assumptions $(i)$, $(ii)$ and $(iii)$, assume
@@
where, for $h,j=2,\dots,N$
@@
First of all, we observe that $\widehat{{\mathbf{A}}}_{t}=UV^{\top}{\mathbf{A}}_{t}$ and the dynamics of $\widehat{{\mathbf{A}}}_{t}$ can be obtained by multypling by $UV^{\top}$ the dynamics of ${\mathbf{A}}_{t}$ as follows:
@@
Recalling that the eigenvalues of $\Phi^{\top}$ different from $\phi^{*}$ are denoted by $\phi_{j}$ for $j=2,\dots,N$, let us set $\alpha_{j}=\phi^{*}-\phi_{j}$. Recall that $\mathcal{R}e(\alpha_{j})>0$ since $\phi^{*}$ is the leading eigenvalue and so $\mathcal{R}e(\phi_{j})<\phi^{*}$. Then, setting an initial time $m_{0}$ such that ${\mathcal{R}e}(\alpha_{j})/(m_{0}+1)<1$ for each $j$, we can write
@@
$\begin{split}C_{k+1,t}&=U_{\boldsymbol{\eta}}A_{k+1,t}V_{\boldsymbol{\eta}}^{\top},\qquad\mbox{and}\\ A_{k+1,t}&=\begin{cases}\prod_{m=k+1}^{t}\left(I-\frac{1}{m+1}S_{Q}\right)=\begin{pmatrix}A^{11}_{k+1,t}&{\boldsymbol{0}}&0\\ 0&a^{22}_{k+1,t}&0\\ A^{31}_{k+1,t}&{\boldsymbol{0}}&A^{33}_{k+1,t}\\ \end{pmatrix}\quad&k=m_{0}-1,\dots,t-1,\\ I\quad&k=t\,.\end{cases}\end{split}$
Notice that the blocks $A^{11}_{k+1,t}$, $A^{31}_{k+1,t}$ and $A^{33}_{k+1,t}$ are all diagonal $(N-1)\times(N-1)$ matrices. Moreover, setting for any $x\in\mathbb{C}$ with ${\mathcal{R}e}(x)/(m_{0}+1)<1$, $p_{m_{0}-1}(x)=1$ and $p_{k}(x)=\prod_{m=m_{0}}^{k}(1-\tfrac{x}{m+1})$ for $k\geq m_{0}$ and $F_{k+1,t}(x)=\frac{p_{t}(x)}{p_{k}(x)}$ for $m_{0}-1\leq k\leq t-1$, we get (see (Aletti et al.,, 2019, Lemma A.5) for similar technical calculations)
@@
To this purpose, setting $B_{{\boldsymbol{\eta}},k+1}=V_{\boldsymbol{\eta}}^{\top}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}\Delta{\mathbf{M}}_{{\boldsymbol{\eta}},k+1}^{\top}V_{\boldsymbol{\eta}}$, $B_{k+1}=V^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}V$, ${\mathbf{b}}_{k+1}=V^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}{\mathbf{v}}$ and $b_{k+1}={\mathbf{v}}^{\top}\Delta{\mathbf{M}}_{k+1}\Delta{\mathbf{M}}_{k+1}^{\top}{\mathbf{v}}$, we observe that
@@
Moreover, with some computations (see (Aletti et al.,, 2020, Lemma A.3) for similar technical calculations) for complex numbers $x$, $y$ such that ${\mathcal{R}e}(x+y)-\phi^{*}>0$ we have